Direction: Perform what is being asked. Follow the steps indicated in the first column of the table to complete the activity. In the second column, write your solution on the spaces provided. Situation: A population of PWD learners in 5 schools in a certain municipality of Quezon Province are () 9, 5, 6, 12, and 15. Suppose that two schools were selected as samples, determine the mean and variance of the sampling distribution of sample mean​

1
Expert's answer
2022-03-20T06:44:17-0400

"\mu(9,5)=7"

"\mu(9,6)=7.5"

"\mu(9,12)=11.5"

"\mu(9,15)=12"

"\mu(5,6)=5.5"

"\mu(5,12)=8.5"

"\mu(5,15)=10"

"\mu(6,12)=9"

"\mu(6,15)=11.5"

"\mu(12,15)=13.5"

"\mu=(7+7.5+11.5+12+5.5+8.5+10+9+11.5+13.5)/10=9.6"

"\sigma=\sqrt{\frac{6.76+4.41+3.61+5.76+16.81+1.54+0.16+0.36+3.61+15.21}{9}}=2.54"